ROTATIONAL SPECTROSCOPY OF 1,1-DIFLUOROPROP-2-YNYL RADICAL, F_{2}C-C≡C-H

\underline{LU KANG}, and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The rotational spectra of 1,1-difluoroprop-2-ynyl radical, F_{2}C-C≡C-H, a fluorine analog of the propargyl radical, have been recorded by Fourier transform microwave (FTMW) spectroscopy. N = 1 - 0 to 4 - 3, a type transitions for K_{a} = 0, 1, and 2, were measured between 6.5 and 26.5 GHz. The spectroscopic constants, including fine and hyperfine constants, were precisely determined. These constants are in good agreement with those predicted by a density functional theory ab initio calculation.